log 15(x+4) = t solve for x
Found 2 solutions by lwsshak3, solver91311:
Answer by lwsshak3(6757)

log 15(x+4) = t solve for x
convert to exponential form: base(10) raised to log of number(t)=number(15(x+4))
10^t=15(x+4)=15x+60
15x=10^t-60
x=(10^t-60)/15

Answer by solver91311(17077)



Not sure whether you mean



or



presumimg base 10.

Whichever way it is, use



to start.


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism